Determine the wavelength of proton that has been accelerated through a pd of 200V.


1
Expert's answer
2021-06-01T13:31:28-0400

Answer

Wavelength of proton can be given as

"\lambda=\frac{h}{\sqrt{2mev}}\\"


"\lambda=\frac{6.62\times10^{-34}}{\sqrt{2(1.67\times10^{-27})(1.6\times10^{-19})(200)}}\\\lambda=202.5nm"